### v1

# Continual Learning: Forget-free Winning Subnetworks for Video Representations

Abstract
Inspired by the Regularized Lottery Ticket Hypothesis (RLTH), which highlights the presence of competitive subnetworks within dense networks for continual learning tasks, we introduce Winning Subnetworks (WSN). This approach utilizes reused weights in dense networks to enhance learning in Task Incremental Learning (TIL) scenarios. To mitigate overfitting in Few-Shot Class Incremental Learning (FSCIL), we have developed WSN variants referred to as the Soft subnetwork (SoftNet). Furthermore, addressing WSN’s limitation of sparse reused weights in Video Incremental Learning (VIL), we propose the Fourier Subneural Operator (FSO). The FSO, operating in Fourier space, adaptively and compactly encodes videos, discovering reusable subnetworks with diverse bandwidths. We have applied FSO’s Fourier representations to various continual learning contexts, including VIL, TIL, and FSCIL. Our extensive experiments across these scenarios demonstrate FSO’s remarkable efficacy in continual learning, significantly enhancing task performance at various convolutional representational levels: it boosts performance in the higher layers for TIL and FSCIL and the lower layers for VIL.
Index Terms:
Continual Learning (CL), Task Incremental Learning (TIL), Video Incremental Learning (VIL), Few-shot Class Incremental Learning (FSCIL), Regularized Lottery Ticket Hypothesis (RLTH), Wining SubNetworks (WSN), Soft-Subnetwork (SoftNet), Fourier Subneural Operator (FSO)
1 Introduction
Continual Learning (CL), also known as Lifelong Learning [82, 66, 96, 24], is a learning paradigm where a series of tasks are learned sequentially. The principle objective of continual learning is to replicate human cognition, characterized by the ability to learn new concepts or skills incrementally throughout one’s lifespan. An optimal continual learning system could facilitate a positive forward and backward transfer, leveraging the knowledge gained from previous tasks to solve new ones while also updating its understanding of previous tasks with the new knowledge. However, building successful continual learning systems is challenging due to the occurrence of catastrophic forgetting or catastrophic interference [55], a phenomenon where the performance of the model on previous tasks significantly deteriorates when it learns new tasks. This can make it challenging to retain the knowledge acquired from previous tasks, ultimately leading to a decline in overall performance. To tackle the catastrophic forgetting problem in continual learning, numerous approaches have been proposed, which can be broadly classified as follows: (1) Regularization-based methods [37, 4, 31, 83, 59] aim to keep the learned information of past tasks during continual training aided by sophisticatedly designed regularization terms, (2) Rehearsal-based methods [64, 65, 5, 6, 67, 68] utilize a set of real or synthesized data from the previous tasks and revisit them, and (3) Architecture-based methods [52, 70, 44, 89, 33, 34] propose to minimize the inter-task interference via newly designed architectural components.
Refer to caption             Refer to caption      Refer to caption Refer to caption
(a) Fixed Backbone (Piggyback, SupSup)       (b) Biased Transfer (PackNet, CLNP)       (c) Selective Reuse Expansion beyond Dense Network (APD)       (d) Selective Reuse Expansion within Network (WSN)
Figure 1: Concept Comparison: (a) Piggyback [52], and SupSup [89] find the optimal binary mask on a fixed backbone network a given task (b) PackNet [53] and CLNP [21] forces the model to reuse all features and weights from previous subnetworks which cause bias in the transfer of knowledge (c) APD [93] selectively reuse and dynamically expand the dense network (d) Our WSN selectively reuse and dynamically expand subnetworks within a dense network. Green edges are reused weights.
Despite the remarkable success of recent works on rehearsal- and architecture-based continual learning, most current methods request external memory as new tasks arrive, making the model difficult to scale to larger and more complex tasks. Rehearsal-based CL requires additional storage to store the replay buffer or generative models, and architecture-based methods leverage additional model capacity to account for new tasks. These trends lead to an essential question: how can we build a memory-efficient CL model that does not exceed the backbone network’s capacity or even requires a much smaller capacity? Several studies have shown that deep neural networks are over-parameterized [18, 23, 43] and thus removing redundant/unnecessary weights can achieve on-par or even better performance than the original dense network. More recently, Lottery Ticket Hypothesis (LTH) [20] demonstrates the existence of sparse subnetworks, named winning tickets, that preserve the performance of a dense network. However, searching for optimal winning tickets during continual learning with iterative pruning methods requires repetitive pruning and retraining for each arriving task, which could be more practical.
To tackle the issues of external replay buffer, capacity, and over-fitting, we suggest a novel CL method, which finds the high-performing Winning SubNetwork referred to as WSN [33] given tasks without the need for retraining and rewinding, as shown in Figure 1 (d). Also, we set previous pruning-based CL approaches [52, 89] (see Figure 1 (a)) to baselines of architectures, which obtain task-specific subnetworks given a pre-trained backbone network. Our WSN incrementally learns model weights and task-adaptive binary masks (the subnetworks) within the neural network. To allow the forward transfer when a model learns on a new task, we reuse the learned subnetwork weights for the previous tasks, however selectively, as opposed to using all the weights [53] (see Figure 1 (b)), that may lead to biased transfer. Further, the WSN eliminates the threat of catastrophic forgetting during continual learning by freezing the subnetwork weights for the previous tasks and does not suffer from the negative transfer, unlike [95] (see Figure 1 (c)), whose subnetwork weights for the previous tasks can be updated when training on the new sessions. Moreover, we observed that subnetworks could overfit to limited task data, potentially reducing their effectiveness on new tasks or datasets, such as in Few-Shot Class Incremental Learning (FSCIL). To address this, we adopted the Regularized Lottery Ticket Hypothesis (RLTH) [34]. This led to the discovery of regularized subnetworks characterized by smoother (soft) masks referred to as Soft-SubNetwork (SoftNet), enhancing their adaptability and performance.
While these conventional architecture-based methods, i.e., WSN and SoftNet, offer solutions to prevent forgetting, they are unsuited for sequential complex Video Continual Learning (VCL) as they reuse a few or all adaptive parameters without finely discretized operations. To enhance neural representation incrementally on complex sequential videos, we propose a novel sequential video compilation method to identify and utilize Lottery tickets (i.e., the weights of complicated oscillatory signals) in frequency space. To achieve this, we define Fourier Subnetwork Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively identifies the most effective Lottery tickets for representing complex periodic signals. In practice, given a backbone and FSO architecture, our method continuously learns to identify input-adaptive subnetwork modules and encode each new video into the corresponding module during sequential training sessions. We extend Fourier representations to various continual learning scenarios, such as Task Incremental Learning (TIL), Video Incremental Learning (VIL), and Few-shot Class Incremental Learning (FSCIL), to demonstrate the effectiveness of FSO.
Our contributions can be summarized as follows:
  • We introduce Fourier Subnetwork Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively, identifies the most effective Lottery tickets for representing complex periodic signals such as sequential video compilation.
  • We have expanded the FSO of Fourier representations to encompass a variety of continual learning scenarios: Video Incremental Learning (VIL), Task Incremental Learning (TIL), and Few-Shot Class Incremental Learning (FSCIL). In our evaluations, the proposed FSO demonstrates superior performance compared to architecture-based continual learning models, such as WSN and SoftNet in TIL, VIL, and FSCIL, respectively, thereby underscoring its exceptional representational power.
2 Related Works
Continual Learning [55, 82, 41, 45], also known as lifelong learning, is the challenge of learning a sequence of tasks continuously while utilizing and preserving previously learned knowledge to improve performance on new tasks. Four major approaches have been proposed to tackle the challenges of continual learning, such as catastrophic forgetting. One such approach is regularization-based methods [37, 4, 31, 83, 59], which aim to reduce catastrophic forgetting by imposing regularization constraints that inhibit changes to the weights or nodes associated with past tasks. Rehearsal-based approaches [64, 5, 6, 67, 17, 79, 68] store small data summaries to the past tasks and replay them during training to retain the acquired knowledge. Some methods in this line of work [73, 2] accommodate the generative model to construct the pseudo-rehearsals for previous tasks. Architecture-based approaches [52, 70, 44, 89, 33, 34] use the additional capacity to expand [91, 95], dynamic representation [92, 74] or isolate [66] model parameters, preserving learned knowledge and preventing forgetting. Both rehearsal and architecture-based methods have shown remarkable efficacy in suppressing catastrophic forgetting but require additional capacity for the task-adaptive parameters [89] or the replay buffers. Prompt-based learning, an emerging transfer learning technique in the field of natural language processing (NLP), harnesses a fixed function to condition the model. This empowers the language model to receive additional instructions for enhancing its performance on downstream tasks. Notably, while L2P [88] stands out as the seminal work that bridges the gap between prompting and continual learning, DualPrompt [87] introduces an innovative approach to affixing complementary prompts to the pre-trained backbone, thereby enabling the acquisition of both task-invariant and task-specific instructions. Additionally, other notable contributions in this field encompass DyTox [19], S-Prompt [86], CODA-P [77], ConStruct-VL [76], ST-Prompt [61], and LGCL [35]. Nonetheless, the central focus of this study is to pinpoint the most optimal winning tickets for convolutional operators in a wide array of continual learning scenarios, both in real and Fourier space.
Architecture-based Continual Learning. While most works aim to increase the performance of continual learners by adding memory, some researchers have focused on building memory and computationally efficient continual learners by using pruning-based constraints. CLNP [21] is one example of a method that selects important neurons for a given task using $\ell_{1}$ regularization to induce sparsity and freezes them to maintain performance. Neurons that are not selected are reinitialized for future task training. Another method, Piggyback [52], trains task-specific binary masks on the weights given a pre-trained model. However, this method does not allow for knowledge transfer among tasks, and its performance highly depends on the quality of the backbone model. HAT [70] proposes task-specific learnable attention vectors to identify important weights per task. The masks are formulated to layerwise cumulative attention vectors during continual learning. A recent method, LL-Tickets [12], shows a sparse subnetwork called lifelong tickets that performs well on all tasks during continual learning. The method searches for more prominent tickets from current ones if the obtained tickets cannot sufficiently learn the new task while maintaining performance on past tasks. However, LL-Tickets require external data to maximize knowledge distillation with learned models for prior tasks, and the ticket expansion process involves retraining and pruning steps. WSN [33] is another method that jointly learns the model and task-adaptive binary masks on subnetworks associated with each task. It attempts to select a small set of weights activated (winning ticket) by reusing weights of the prior subnetworks in Task Incremental Learning (TIL). We also investigate the performance of Soft-Subnetworks (SoftNet) [34]. We select sub-networks (winning tickets) to obtain base session knowledge and optimize the regularized sub-network parameters for new sessions in Few-Shot Class Incremental Learning (FSCIL) settings.
Neural Implicit Representation (NIR) [56] are neural network architectures for parameterizing continuous, differentiable signals. Based on coordinate information, they provide a way to represent complex, high-dimensional data with a small set of learnable parameters that can be used for various tasks such as image reconstruction [75, 80], shape regression [13, 60], and 3D view synthesis  [58, 69]. Instead of using coordinate-based methods, NeRV [10] proposes an image-wise implicit representation that takes frame indices as inputs, enabling fast and accurate video compression. NeRV has inspired further improvements in video regression by CNeRV [8], DNeRV [25], E-NeRV [46], and NIRVANA [51], and HNeRV [9]. A few recent works have explored video continual learning (VCL) scenarios for the NIR. To tackle non-physical environments, Continual Predictive Learning (CPL) [7] learns a mixture world model via predictive experience replay and performs test-time adaptation using non-parametric task inference. PIVOT [85] leverages the past knowledge present in pre-trained models from the image domain to reduce the number of trainable parameters and mitigate forgetting. CPL needs memory to replay, while PIVOT needs pre-training and fine-tuning steps. In contrast, along with the conventional progressive training techniques [66, 16], we use WSN as baselines, which utilizes the Lottery Ticket Hypothesis (LTH) to identify an adaptive substructure within the dense networks that are tailored to the specific video input index. However, WSN is inappropriate for sequential complex video compilation since it reuses a few adaptive but sparse learnable parameters. To overcome the weakness of WSN, We proposed a novel Fourier Subneural Operator (FSO) for representing complex video in Fourier space [47, 48, 38, 84]. We have expanded the FSO of Fourier representations to encompass a variety of continual learning architectures and scenarios to validate its effectivness.
3 Winning SubNetworks in Fourier Space
In this section, we present our pruning-based continual learning methods, Winning SubNetworks (WSN, see Figure 2[33]. The WSN searches for the task-adaptive winning tickets and updates only the weights that have not been trained on the previous tasks, as shown in Figure 2. After training on each session, the subnetwork parameters of the model are frozen to ensure that the proposed method is inherently immune to catastrophic forgetting. Moreover, we illustrate Soft-Winning SubNetworks (SoftNet) [34], proposed to address the issues of forgetting previous sessions and overfitting a few samples of new sessions. These conventional architecture-based methods, i.e., WSN and SoftNet, offer solutions to prevent forgetting, however, they are unsuited for sequential complex Video Incremental Learning (VIL see Figure 3) as they reuse a few or all adaptive parameters without finely discretized operations. To enhance neural representation incrementally on complex sequential videos, we introduce Fourier Subnetwork Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively identifies the most effective Lottery tickets for representing complex periodic signals. In practice, given a backbone and FSO architecture, our method continuously learns to identify input-adaptive subnetwork modules and encode each new video into the corresponding module during sequential training sessions. We extend Fourier representations to various continual learning scenarios, such as TIL, VIL, and FSCIL, to demonstrate the effectiveness of FSO.
Refer to caption           Refer to caption           Refer to caption           Refer to caption
      (a) selected weights at prior task (b) forward pass at current task (c) backward pass at current task (d) selected weights at current task
Figure 2: An illustration of Winning SubNetworks (WSN): (a) The top-cweights $\hat{\bm{\theta}}_{s-1}$ at prior task are obtained, (b) In the forward pass of a new task, WSN reuses weights selected from prior tasks, (c) In the backward pass, WSN updates only non-used weights, and (d) after several iterations of (b) and (c), we acquire again the top-cweights $\hat{\bm{\theta}}_{s}$ including subsets of reused weights (green) for the new task.
Problem Statement. Consider a supervised learning setup where $S$ sessions or tasks arrive to a learner sequentially. We denote that $\mathcal{D}_{s}=\{\textbf{x}_{i,s},y_{i,s}\}_{i=1}^{n_{s}}$ is the dataset of session $s$, composed of $n_{s}$ pairs of raw instances and corresponding labels. We assume a neural network $f(\cdot;\bm{\theta})$, parameterized by the model weights $\bm{\theta}$ and standard continual learning scenario aims to learn a sequence of sessions by solving the following optimization procedure at each step $s$:
$\bm{\theta}^{\ast}=\operatorname*{minimize}_{\bm{\theta}}\frac{1}{n_{s}}\sum^{n_{s}}_{i=1}\mathcal{L}(f(\textbf{x}_{i,s};\bm{\theta}),y_{i,s}),$ (1)
where $\mathcal{L}(\cdot,\cdot)$ is a classification objective loss such as cross-entropy loss. $\mathcal{D}_{s}$ for session $s$ is only accessible when learning session $s$.
Continual learners frequently use over-parameterized deep neural networks to ensure enough capacity for learning future tasks. This approach often leads to the discovery of subnetworks that perform as well as or better than the dense network. Given the neural network parameters $\bm{\theta}$, the binary attention mask $\bm{m}^{*}_{s}$ that describes the optimal subnetwork for session $s$ such that $|\bm{m}^{*}_{s}|$ is less than the model capacity $c$ follows as:
$\begin{split}\bm{m}^{*}_{s}&=\underset{\bm{m}_{s}\in\{0,1\}^{|\bm{\theta}|}}{\operatorname*{minimize}}\frac{1}{n_{s}}\sum^{n_{s}}_{i=1}\mathcal{L}\big{(}f(\textbf{x}_{i,s};\bm{\theta}\odot\bm{m}_{s}),y_{i,s}\big{)}-\mathcal{J}\\ &\quad\quad\quad\quad\quad\quad\text{subject to\leavevmode\nobreak\ }|\bm{m}^{*}_{s}|\leq c,\end{split}$ (2)
[…]
3.1 Fourier Subnueral Operator (FSO)
Conventional continual learner (i.e., WSN) only uses a few learnable parameters in convolutional operations to represent complex sequential image streams in Video Incremental Learning. To capture more parameter-efficient and forget-free video representations (i.e., Neural Implicit Representation (NIR), see Figure 3), the NIR model requires fine discretization and video-specific sub-parameters. This motivation leads us to propose a novel subnetwork operator in Fourier space, which provides it with various bandwidths. Following the previous definition of Fourier convolutional operator [47], we adapt and redefine this definition to better fit the needs of the NIR framework. We use the symbol $\mathcal{F}$ to represent the Fourier transform of a function $f$, which maps from an embedding space of dimension $d_{\bm{e}}=1\times 160$ to a frame size denoted as $d_{\bm{v}}$. The inverse of this transformation is represented by $\mathcal{F}^{-1}$. In this context, we introduce our Fourier-integral Subneural Operator (FSO), symbolized as $\mathcal{K}$, which is tailored to enhance the capabilities of our NIR system:
$\left(\mathcal{K}(\phi)\bm{e}_{s,t}\right)(\bm{v}_{t}^{s})=\mathcal{F}^{-1}(R_{\bm{\phi}}\cdot(\mathcal{F}\bm{e}_{s,t}))(\bm{v}_{t}^{s}),$ (3)
where $R_{\bm{\phi}}$ is the Fourier transform of a periodic subnetwork function which is parameterized by its subnetwork’s parameters of real $(\bm{\theta}^{real}\odot\bm{m}_{s}^{real})$ and imaginary $(\bm{\theta}^{imag}\odot\bm{m}_{s}^{imag})$. We thus parameterize $R_{\bm{\phi}}$ separately as complex-valued tensors of real and imaginary $\bm{\phi}_{FSO}\in\{\bm{\theta}^{real},\bm{\theta}^{imag}\}$. One key aspect of the FSO is that its parameters grow with the depth of the layer and the input/output size. However, through careful layer-wise inspection and adjustments for sparsity, we can find a balance that allows the FSO to describe neural implicit representations efficiently. In the experimental section, we will showcase the most efficient FSO structure and its performance. Figure 3 shows one possible structure of a single FSO. We describe the optimization in the following section.
Figure 3: Forget-free Neural Implicit Representation with Fourier Subneural Operator (FSO) for Video Incremental Learning: Image-wise neural implicit representation taking frame and video (session $s$) indices as input and using a sparse Stems + NeRV Blocks with Fourier Subneural Operator (FSO) to output the whole image through multi-heads $H_{N}$. We denote frozen, reused, and trainable parameters in training at session 2. Note that each video representation is colored. In inference, we only need indices of session $s$ and frame $t$ and session mask (subnetwork).
For Task/Class Incremental Learnings, Convolutional Neural Networks (CNNs) take a convolutional operation, followed by a pooling operation. These iterative operations of CNNs represent more abstract features at higher layer output levels and lose global contextual representations. To compensate for low-contextual representations, we add an FSO to CNN architecture as shown in Figure 4. The lower layer’s output $\bm{x}^{l-1}$ is merged into the $l$th layer residual block through the FSO to acquire spatially ensembling features $\bm{x}^{l}$. The FSO also provides additional parameters to push the residual to zero. We show the differences between ensembling features (WSN+FSO) and single features (WSN) represented by residual blocks as shown in Figure 5: WSN+FSO provides lower variances of feature maps and higher frequency components than WSN. In the following experimental settings, we investigate various CNN architectures with FSO.
Figure 4: Residual Blocks (ResBlocks) with Fourier Subnerual Operator (FSO).
Refer to caption Refer to caption
(a) feature map variance (b) high-freq. of feature map
Figure 5: The comparisons of WSN of 4 Conv (blue area) and 3 FC (gray area) with FSO (white area) in terms of Feature variances and high-frequency components: the (a) offers the variance of the feature map and the (b) provides $\Delta\log$ amplitudes at high-frequency ($1.0\pi$).
3.2 SubNetworks with FSO
3.2.1 Winning SubNetworks (WSN) with FSO
Let each weight $\bm{\theta}_{\ast}=\{\bm{\theta},\bm{\phi}_{FSO}\}$ be associated with a learnable parameter we call weight score $\bm{\rho}_{\ast}=\{\bm{\rho},\bm{\rho}_{FSO}\}$, which numerically determines the importance of the weight associated with it; that is, a weight with a higher weight score is seen as more important. We find a sparse subnetwork $\hat{\bm{\theta}}_{s}$ of the neural network and assign it as a solver of the current session $s$. We use subnetworks instead of the dense network as solvers for two reasons: (1) Lottery Ticket Hypothesis [20] shows the existence of a subnetwork that performs as well as the whole network, and (2) subnetwork requires less capacity than dense networks, and therefore it inherently reduces the size of the expansion of the solver.
Motivated by such benefits, we propose a novel Winning SubNetworks (WSN [^1] WSN code is available at https://github.com/ihaeyong/WSN.git), which is the joint-training method for continual learning that trains on session - while selecting an important subnetwork given the session $s$ as shown in Fig. 2. The illustration of WSN explains step-by-step how to acquire binary weights within a dense network. We find $\hat{\bm{\theta}}_{t}$ by selecting the $c$weights with the highest weight scores $\bm{\rho}_{\ast}$, where $c$ is the target layerwise capacity ratio in %. A task-dependent binary weight represents the selection of weights $\bm{m}_{s}$ where a value of $1$ denotes that the weight is selected during the forward pass and 00 otherwise. Formally, $\mathbf{m}_{s}$ is obtained by applying a indicator function $\mathbbm{1}_{c}$ on $\bm{\rho}$ where $\mathbbm{1}_{c}(\rho)=1$ if $\rho_{\ast}$ belongs to top-$c\%$ scores and 00 otherwise. Therefore, the subnetwork $\hat{\bm{\theta}}_{s}$ for session $s$ is obtained by $\hat{\bm{\theta}}_{s}=\bm{\theta}_{\ast}\odot\mathbf{m}_{s}$.
3.2.2 Soft-Subnetworks (SoftNet) with FSO
Several works have addressed overfitting issues in continual learning from different perspectives, including NCM [30], BiC [90], OCS [94], and FSLL [54]. To mitigate the overfitting issue in subnetworks, we use a simple yet efficient method named SoftNet proposed by [34]. The following new paradigm, referred to as Regularized Lottery Ticket Hypothesis [34] which is inspired by the Lottery Ticket Hypothesis [20] has become the cornerstone of SoftNet:
Regularized Lottery Ticket Hypothesis (RLTH). A randomly-initialized dense neural network contains a regularized subnetwork that can retain the prior class knowledge while providing room to learn the new class knowledge through isolated training of the subnetwork.
[…]
3.3 Optimization for TIL, VIL, and FSCIL
3.3.1 Winning SubNetworks (WSN) for TIL
To jointly learn the model weights and task-adaptive binary masks of subnetworks associated with each session, given an objective $\mathcal{L}(\cdot)$, i.e., cross-entropy loss, we optimize $\bm{\theta}_{\ast}$ and $\bm{\rho}_{\ast}$ with:
$\operatorname*{minimize}_{\bm{\theta}_{\ast},\bm{\rho}_{\ast}}\mathcal{L}(\bm{\theta}_{\ast}\odot\bm{m}_{s};\mathcal{D}_{s}).$ (4)
However, this vanilla optimization procedure presents two problems: (1) updating all $\bm{\theta}_{\ast}$ when training for new sessions will cause interference to the weights allocated for previous sessions, and (2) the indicator function always has a gradient value of 00; therefore, updating the weight scores $\bm{\rho}_{\ast}$ with its loss gradient is not possible. To solve the first problem, we selectively update the weights by allowing updates only on the weights not selected in the previous sessions. To do that, we use an accumulate binary mask $\bm{M}_{s-1}=\operatorname*{\vee}_{i=1}^{s-1}\bm{m}_{i}$ when learning session $s$, then for an optimizer with learning rate $\eta,$ the $\bm{\theta}_{\ast}$ is updated as follows:
$\bm{\theta}_{\ast}\leftarrow\bm{\theta}_{\ast}-\eta\left(\frac{\partial\mathcal{L}}{\partial\bm{\theta}_{\ast}}\odot(\bm{1}-\bm{M}_{s-1})\right),$ (5)
effectively freezing the weights of the subnetworks selected for the previous sessions. To solve the second problem, we use Straight-through Estimator [29, 3, 63] in the backward pass since $\bm{m}_{s}$ is obtained by top-$c\%$ scores. Specifically, we ignore the derivatives of the indicator function and update the weight score as follows:
$\quad\bm{\rho}_{\ast}\leftarrow\bm{\rho}_{\ast}-\eta\left(\frac{\partial\mathcal{L}}{\partial\bm{\rho}_{\ast}}\right).$ (6)
Our WSN optimizing procedure is summarized in Algorithm 1.
Algorithm 1 Winning SubNetworks (WSN) for TIL and VIL.
0:  $\{\mathcal{D}_{s}\}_{s=1}^{\mathcal{S}}$, model weights $\bm{\theta}_{\ast}=\{\bm{\theta},\bm{\phi}_{FSO}\}$, score weights $\bm{\rho}_{\ast}=\{\bm{\rho},\bm{\rho}_{FSO}\}$, binary mask $\mathbf{M}_{0}=\{\bm{0}^{|\bm{\theta}|},\bm{0}^{|\bm{\theta}_{FSO}|}\}$, layer-wise capacity $c\%$.
1:  Randomly initialize $\bm{\theta}_{\ast}$ and $\bm{\rho}_{\ast}$.
2:  for task $s=1,\cdots,\mathcal{S}$ do
3:    for batch $\bm{b}_{s}\sim\mathcal{D}_{s}$ do
4:       Obtain mask $\bm{m}_{s}$ of the top-$c\%$ scores $\bm{\rho}_{\ast}$
5:       Compute $\mathcal{L}\left(\bm{\theta}_{\ast}\odot\bm{m}_{s};\bm{b}_{s}\right)$
6:       $\bm{\theta}_{\ast}\leftarrow\bm{\theta}_{\ast}-\eta\left(\frac{\partial\mathcal{L}}{\partial\bm{\theta}_{\ast}}\odot(\bm{1}-\bm{M}_{s-1})\right)$triangleright Weight update
7:       $\bm{\rho}_{\ast}\leftarrow\bm{\rho}_{\ast}-\eta(\frac{\partial\mathcal{L}}{\partial\bm{\rho}_{\ast}})$triangleright Weight score update
8:    end for
[…]
10:  end for
3.3.2 Winning SubNetworks (WSN) for VIL
Let a video at $s_{th}$ session $\bm{V}_{s}=\{\bm{v}^{s}_{t}\}^{T_{s}}_{t=1}\in\mathbb{R}^{T_{s}\times H\times W\times 3}$ be represented by a function with the trainable parameter $\bm{\theta}_{\ast}$, $f_{\bm{\theta}_{\ast}}:\mathbb{R}\rightarrow\mathbb{R}^{H\times W\times 3}$, during Video Incremental Learning (VIL), where $T_{s}$ denotes the number of frames in a video at session $s$, and $s\in\{1\dots,|S|\}$. Given a session and frame index $s$ and $t$, respectively, the neural implicit representation aims to predict a corresponding RGB image $\bm{v}^{s}_{t}\in\mathbb{R}^{H\times W\times 3}$ by fitting an encoding function to a neural network: $\bm{v}^{s}_{t}=f_{\bm{\theta}_{\ast}}([s;t],H_{s})$ where $H_{s}$ is $s_{th}$ head. For the sake of simplicity, we omit $H_{s}$ in the following equations. Let’s consider a real-world learning scenario in which $|\mathcal{S}|=N$ or more sessions arrive in the model sequentially. We denote that $\mathcal{D}_{s}=\{\bm{e}_{s,t},\bm{v}_{s,t}\}_{t=1}^{T_{s}}$ is the dataset of session $s$, composed of $T_{s}$ pairs of raw embeddings $\bm{e}_{s,t}=\left[\bm{e}_{s};\bm{e}_{t}\right]\in\mathbb{R}^{1\times 160}$ and corresponding frames $\bm{v}^{s}_{t}$. Here, we assume that $\mathcal{D}_{s}$ for session $s$ is only accessible when learning session $s$ due to the limited hardware memory and privacy-preserving issues, and session identity is given in the training and testing stages. The primary training objective in this sequence of $N$ video sessions is to minimize the following optimization problem:
$\operatorname*{minimize}_{\bm{\theta}_{\ast},\bm{\rho}_{\ast}}\frac{1}{N}\frac{1}{T_{s}}\sum^{N}_{s=1}\sum^{T_{s}}_{t=1}\mathcal{L}(f(\bm{e}_{s,t};\bm{\theta}_{\ast}\odot\bm{m}_{s}),\bm{v}^{s}_{t}),$ (7)
where the loss function $\mathcal{L}(\bm{v}^{s}_{t})$ is composed of $\ell_{1}$ loss and SSIM loss. The former minimizes the pixel-wise RGB gap with the original input frames evenly, and the latter maximizes the similarity between the two entire frames based on luminance, contrast, and structure, as follows:
$\mathcal{L}(\bm{V}_{s})={1\over T_{s}}\sum_{t=1}^{T_{s}}\alpha||\bm{v}^{s}_{t}-\hat{\bm{v}}^{s}_{t}||_{1}+(1-\alpha)(1-\text{SSIM}(\bm{v}^{s}_{t},\hat{\bm{v}}^{s}_{t})),$ (8)
where $\hat{\bm{v}}^{s}_{t}$ is the output generated by the model $f$. For all experiments, we set the hyperparameter $\alpha$ to $0.7$, and we adapt PixelShuffle [72] for session and time positional embedding.
3.3.3 Soft-SubNetworks (SoftNet) for FSCIL
Similar to WSN’s optimization discussed in Section 3.3, let each weight $\bm{\theta}_{\ast}$ be associated with a learnable parameter we call weight score $\bm{\rho}_{\ast}$, which numerically determines the importance of the associated weight. In the optimization process for FSCIL, however, we consider two main problems: (1) Catastrophic forgetting: updating all $\bm{\theta}_{\ast}\odot\bm{m}_{s-1}$ when training for new sessions will cause interference with the weights allocated for previous sessions; thus, we need to freeze all previously learned parameters $\bm{\theta}_{\ast}\odot\bm{m}_{s-1}$; (2) Overfitting: the subnetwork also encounters overfitting issues when training an incremental session on a few samples, as such, we need to update a few parameters irrelevant to previous session knowledge., i.e., $\bm{\theta}_{\ast}\odot(\bm{1}-\bm{m}_{s-1})$.
To acquire the optimal subnetworks that alleviate the two issues, we define a soft-subnetwork by dividing the dense neural network into two parts-one is the major subnetwork $\bm{m}_{\text{major}}$, and another is the minor subnetwork $\bm{m}_{\text{minor}}$. The defined Soft-SubNetwork (SoftNet) follows as:
$\bm{m}_{\text{soft}}=\bm{m}_{\text{major}}\oplus\bm{m}_{\text{minor}},$ (9)
where $\bm{m}_{\text{major}}$ is a binary mask and $\bm{m}_{\text{minor}}\sim U(0,1)$ and $\oplus$ represents an element-wise summation. As such, a soft-mask is given as $\bm{m}_{s}^{\ast}\in[0,1]^{|\bm{\theta}_{\ast}|}$. In the all-experimental FSCIL setting, $\bm{m}_{\text{major}}$ maintains the base session knowledge $s=1$ while $\bm{m}_{\text{minor}}$ acquires the novel session knowledge $s\geq 2$. Then, with base session learning rate $\alpha,$ the $\bm{\theta}_{\ast}$ is updated as follows: $\bm{\theta}_{\ast}\leftarrow\bm{\theta}_{\ast}-\alpha\left(\frac{\partial\mathcal{L}}{\partial\bm{\theta}_{\ast}}\odot\bm{m}_{\text{soft}}\right)$ effectively regularize the weights of the subnetworks for incremental learning. Our Soft-subnetwork optimizing procedure is summarized in Algorithm 2. Once a single soft-subnetwork $\bm{m}_{\text{soft}}$ is obtained in the base session, then we use the soft-subnetwork for the entire new sessions without updating.
Algorithm 2 Soft-Subnetworks (SoftNet) for FSCIL.
0:  $\{\mathcal{D}^{t}\}_{t=1}^{\mathcal{T}}$, model weights $\bm{\theta}_{\ast}=\{\bm{\theta},\bm{\phi}_{FSO}\}$, and score weights $\bm{\rho}_{\ast}=\{\bm{\rho},\bm{\rho}_{FSO}\}$, layer-wise capacity $c$
1:  // Training over base classes $s=1$
2:  Randomly initialize $\bm{\theta}_{\ast}$ and $\bm{\rho}_{\ast}$.
3:  for epoch $e=1,2,\cdots$ do
4:    Obtain softmask $\bm{m}_{\text{soft}}$ of $\bm{m}_{major}$ and $\bm{m}_{minor}\sim U(0,1)$ at each layer
5:    for batch $\bm{b}_{s}\sim\mathcal{D}^{s}$ do
6:       Compute $\mathcal{L}_{base}\left(\bm{\theta}_{\ast}\odot\bm{m}_{\text{soft}};\bm{b}_{s}\right)$ by Equation 4
7:       $\bm{\theta}_{\ast}\leftarrow\bm{\theta}_{\ast}-\alpha\left(\frac{\partial\mathcal{L}}{\partial\bm{\theta}_{\ast}}\odot\bm{m}_{\text{soft}}\right)$
8:       $\bm{\rho}_{\ast}\leftarrow\bm{\rho}_{\ast}-\alpha\left(\frac{\partial\mathcal{L}}{\partial\bm{\rho}_{\ast}}\odot\bm{m}_{\text{soft}}\right)$
9:    end for
10:  end for
11:  // Incremental learning $s\geq 2$
12:  Combine the training data $\mathcal{D}^{s}$
13:     and the exemplars saved in previous few-shot sessions
14:  for epoch $e=1,2,\cdots$ do
15:    for batch $\bm{b}_{s}\sim\mathcal{D}^{s}$ do
16:       Compute $\mathcal{L}_{m}\left(\bm{\theta}_{\ast}\odot\bm{m}_{\text{soft}};\bm{b}_{s}\right)$ by Equation 10
17:       $\bm{\theta}_{\ast}\leftarrow\bm{\theta}_{\ast}-\beta\left(\frac{\partial\mathcal{L}}{\partial\bm{\theta}_{\ast}}\odot\bm{m}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}minor}}\right)$
18:    end for
19:  end for
19:  model parameters $\bm{\theta}_{\ast}$, $\bm{\rho}_{\ast}$, and $\bm{m}_{\text{soft}}$.
Base Training $(s=1)$. In the base learning session, we optimize the soft-subnetwork parameter $\bm{\theta}_{\ast}$ (including a fully connected layer as a classifier) and weight score $\bm{\rho}_{\ast}$ with cross-entropy loss jointly using the training examples of $\mathcal{D}^{1}$.
Incremental Training $(s\geq 2)$. In the incremental few-shot learning sessions $(s\geq 2)$, leveraged by $\bm{\theta}_{\ast}\odot\bm{m}_{\text{soft}}$, we fine-tune few minor parameters $\bm{\theta}_{\ast}\odot\bm{m}_{\text{minor}}$ of the soft-subnetwork to learn new classes. Since $\bm{m}_{\text{minor}}<\bm{1}$, the soft-subnetwork alleviates the overfitting of a few samples. Furthermore, instead of Euclidean distance [71], we employ a metric-based classification algorithm with cosine distance to finetune the few selected parameters. In some cases, Euclidean distance fails to give the real distances between representations, especially when two points with the same distance from prototypes do not fall in the same class. In contrast, representations with a low cosine distance are located in the same direction from the origin, providing a normalized informative measurement. We define the loss function as:
$\begin{split}\mathcal{L}_{m}(\bm{x},y;\bm{\theta}_{\ast}\odot\bm{m}_{soft})=\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ &\\ -\sum_{\bm{x},y\in\mathcal{D}}\sum_{o\in\mathcal{O}}\mathbbm{1}(y=o)\log\left(\frac{e^{-d(\bm{p}_{o},f(\bm{x};\;\bm{\theta}_{\ast}\odot\bm{m}_{soft}))}}{\sum_{o_{k}\in\mathcal{O}}e^{-d(\bm{p}_{o_{k}},f(\bm{x};\;\bm{\theta}_{\ast}\odot\bm{m}_{soft}))}}\right)\end{split}$ (10)
where $d\left(\cdot,\cdot\right)$ denotes cosine distance, $\bm{p}_{o}$ is the prototype of class $o$, $\mathcal{O}=\bigcup_{i=1}^{s}\mathcal{O}^{i}$ refers to all encountered classes, and $\mathcal{D}=\mathcal{D}^{s}\bigcup\mathcal{P}$ denotes the union of the current training data $\mathcal{D}^{s}$ and the exemplar set $\mathcal{P}=\left\{\bm{p}_{2}\cdots,\bm{p}_{s-1}\right\}$, where $\mathcal{P}_{s_{e}}\left(2\leq s_{e}<s\right)$ is the set of saved exemplars in session $s_{e}$. Note that the prototypes of new classes are computed by $\bm{p}_{o}=\frac{1}{N_{o}}\sum_{i}\mathbbm{1}(y_{i}=o)f(\bm{x}_{i};\bm{\theta}_{\ast}\odot\bm{m}_{soft})$ and those of base classes are saved in the base session, and $N_{o}$ denotes the number of the training images of class $o$. We also save the prototypes of all classes in $\mathcal{O}^{s}$ for later evaluation.
Inference for Incremental Soft-Subnetwork. In each session, the inference is also conducted by a simple nearest class mean (NCM) classification algorithm [57, 71] for fair comparisons. Specifically, all the training and test samples are mapped to the embedding space of the feature extractor $f$, and Euclidean distance $d_{u}(\cdot,\cdot)$ is used to measure their similarity. The classifier gives the $k$th prototype index $o_{k}^{\ast}=\arg\min_{o\in\mathcal{O}}d_{u}(f(\bm{x};\bm{\theta}_{\ast}\odot\bm{m}_{soft}),\bm{p}_{o})$ as output.
4 Experiments
We validate our method on several benchmark datasets against relevant continual learning baselines on Task-Incremental Learning (TIL), Video Incremental Learning (VIL), and Few-shot Class Incremental Learning (FSCIL).
4.1 Task-incremental Learning (TIL)
Datasets and architectures. We use three different popular sequential datasets for CL problems with three different neural network architectures as follows: 1) CIFAR-100 Split [39]: A visual object dataset constructed by randomly dividing 100 classes of CIFAR-100 into ten tasks with ten classes per task. 2) CIFAR-100 Superclass: We follow the setting from [93] that divides CIFAR-100 dataset into 20 tasks according to the 20 superclasses, and each superclass contains five different but semantically related classes. 3) TinyImageNet [78]: A variant of ImageNet [40] containing 40 of 5-way classification tasks with the image size by $64\times 64\times 3$.
We use variants of LeNet [42] for the experiments on CIFAR-100 Superclass experiments, and a modified version of AlexNet similar to [70, 67] for the CIFAR-100 Split dataset. For TinyImageNet, we also use the same network architecture as [22, 17], which consists of 4 Conv layers and 3 fully connected layers. We set WSN as the architecture-based baselines, which jointly train and find task-adaptive subnetworks of novel/prior parameters for continual learning. WSN+FSO follows the Residual Blocks as stated in Figure 4 in all three architectures. The FSO’s ensemble structures follow in terms of three architectures:
  • In the LeNet, the Conv. layer’s output $\bm{x}^{l=1}$ is merged into the $l=2$th Conv. layer output through the FSO to acquire spatially ensembling features $\bm{x}^{l=2}$.
  • In the AlexNet, the Conv. layer’s output $\bm{x}^{l=2}$ is merged into the $l=3$th Conv. layer’s output through the FSO to acquire spatially ensembling features $\bm{x}^{l=3}$.
  • In the TinyImageNet, the Conv. layer’s output $\bm{x}^{l=2}$ is merged into the $l=3$th Conv. layer’s output through the FSO to acquire spatially ensembling features $\bm{x}^{l=3}$.
Experimental settings. As we directly implement our method from the official code of [67], we provide the values for HAT and GPM reported in [67]. For Omniglot Rotation and Split CIFAR-100 Superclass, we deploy the proposed architecture in multi-head settings with hyperparameters as reported in [93]. All our experiments run on a single-GPU setup of NVIDIA V100.
Performance metrics. We evaluate all methods based on the following four metrics:
  1. 1.
    Accuracy (ACC) measures the average of the final classification accuracy on all tasks: $\mathrm{ACC}=\frac{1}{T}\sum_{i=1}^{T}A_{T,i}$, where $A_{T,i}$ is the test accuracy for task $i$ after training on task $T$.
  2. 2.
    Backward Transfer (BWT) measures the forgetting during continual learning. Negative BWT means that learning new tasks causes the forgetting of past tasks: $\mathrm{BWT}=\frac{1}{T-1}\sum_{i=1}^{T-1}A_{T,i}-A_{i,i}$.
Baselines. We compare our WSN with strong CL baselines; regularization-based methods: HAT [70] and EWC [37], rehearsal-based methods: GPM [67], and a pruning-based method: PackNet [53] and SupSup [89]. PackNet and SupSup is set to the baseline to show the effectiveness of re-used weights. We also compare with a naive sequential training strategy, referred to as FINETUNE. Multitask Learning (MTL) and Single-task Learning (STL) are not a CL method. MTL trains on multiple tasks simultaneously, and STL trains on single tasks independently.
4.2 Video incremental Learning (VIL)
We validate our method on video benchmark datasets against continual learning baselines on Video Incremental Learning (VIL). We consider continual video representation learning with a multi-head configuration (session id, i.e., $s$ is given in training and inference) for all experiments in the paper. We follow the experimental setups in NeRV [10] and HNeRV [9].
Datasets and architectures. We conducted an extended experiment on the UVG of 8/17 video sessions. The category index and order in UVG8 (1.bunny, 2.beauty, 3.bosphorus, 4.bee, 5.jockey, 6.setgo, 7.shake, 8.yacht) and UVG17 (1.bunny, 2.city, 3.beauty, 4.focus, 5.bosphorus, 6.kids, 7.bee, 8.pan, 9.jockey, 10.lips, 11.setgo, 12.race, 13.shake, 14.river, 15.yacht, 16.sunbath, 17.twilight). We employ NeRV as our baseline architecture and follow its details for a fair comparison. After the positional encoding, we apply 2 sparse MLP layers on the output of the positional encoding layer, followed by five sparse NeRV blocks with upscale factors of 5, 2, 2, 2, 2. These sparse NeRV blocks decode 1280$\times$720 frames from the 16$\times$9 feature map obtained after the sparse MLP layers. For the upscaling method in the sparse NeRV blocks, we also adopt PixelShuffle [72]. Fourier Subneural Operator (FSO) is used at the NeRV2 or NeRV3 layer, denoted as $f$-NeRV2 and $f$-NeRV3333The $f$-NeRV3 code is available at https://github.com/ihaeyong/PFNR.git). The positional encoding for the video index $s$ and frame index $t$ is as follows:
$\begin{split}\bm{\Gamma}(s,t)=&[\leavevmode\nobreak\ \sin(b^{0}\pi s),\cos(b^{0}\pi s),\cdots,\sin(b^{l-1}\pi s),\cos(b^{l-1}\pi s),\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \sin(b^{0}\pi t),\cos(b^{0}\pi t),\cdots,\sin(b^{l-1}\pi t),\cos(b^{l-1}\pi t)\leavevmode\nobreak\ ],\end{split}$ (11)
where the hyperparameters are set to $b=1.25$ and $l=80$ such that $\bm{\Gamma}(s,t)\in\mathbb{R}^{1\times 160}$. As differences from the previous NeRV model, the first layer of the MLP has its input size expanded from 80 to 160 to incorporate both frame and video indices, and distinct head layers after the NeRV block are utilized for each video. For the loss objective in Equation 8, $\alpha$ is set to $0.7$. We evaluate the video quality, average video session quality, and backward transfer with two metrics: PSNR.
Baselines. To show the effectiveness, we compare our WSN+FSO with strong CL baselines: Single-Task Learning (STL), which trains on single tasks independently, EWC [37], which is a regularized baseline, iCaRL [64], and ESMER [68] which are current strong rehearsal-based baseline, WSN [33] which is a current strong architecture-based baseline, and Multi-Task Learning (MTL) which trains on multiple video sessions simultaneously, showing the upper-bound of WSN. Except for STL, all models are trained and evaluated on multi-head settings where a video session and time $(s,t)$ indices are provided.
Training. In all experiments, we follow the same experimental settings as NeRV [9] and HNeRV [9] for fair comparisons. We train WSN+FSO, NeRV (STL), and MTL using Adam optimizer with a learning rate 5e-4. For the ablation study on UVG17, we use a cosine annealing learning rate schedule [50], batch size of 1, training epochs of 150, and warmup epochs of 30 unless otherwise denoted.
VIL’s performance metrics of PSNR We evaluate all methods based on the following continual learning metrics:
  1. 1.
    Average PSNR measures the average of the final performances on all video sessions: $\mathrm{PSNR}=\frac{1}{N}\sum_{s=1}^{N}A_{N,s}$, where $A_{N,s}$ is the test PSNR for session $s$ after training on the final video session $S$.
  2. 2.
    Backward Transfer (BWT) of PSNR measures the video representation forgetting during continual learning. Negative BWT means that learning new video sessions causes the video representation forgetting of past sessions: $\mathrm{BWT}=\frac{1}{N-1}\sum_{s=1}^{N-1}A_{N,s}-A_{s,s}$.
4.3 Few-shot Class Incremental Learning (FSCIL)
We introduce experimental setups - Few-Shot Class Incremental Learning (FSCIL) settings to provide soft-subnetworks’ effectiveness. We empirically evaluate and compare our soft subnetworks with state-of-the-art methods and vanilla subnetworks in the following subsections.
Datasets. To validate the effectiveness of the soft subnetwork, we follow the standard FSCIL experimental setting. We randomly selected 60 classes as the base and 40 as new classes for CIFAR-100 and miniImageNet. In each incremental learning session, we construct 5-way 5-shot tasks by randomly picking five classes and sampling five training examples for each class; we set the first 100 classes of CUB-200-2011 as base classes and the remaining 100 classes as new categories split into 10 novel sessions (i.e., a 10-way 5-shot).
Baselines. We mainly compare our SoftNet [34] with architecture-based methods for FSCIL: FSLL [54] that selects important parameters for each session, and HardNet, representing a binary subnetwork. Furthermore, we compare other FSCIL methods such as iCaRL [64], Rebalance [30], TOPIC [81], IDLVQ-C [11], and F2M [71]. Fourier Subneural Operator (FSO) is used at the 3th residual blocks of ResNet18. The $3$th residual block’s output $\bm{x}^{l=3}$ is merged into the $l=4$th residual block through the FSO to acquire spatially ensembling features $\bm{x}^{l=4}$. We also include a joint training method [71] that uses all previously seen data, including the base and the following few-shot tasks for training as a reference. Furthermore, we fix the classifier re-training method (cRT) [32] for long-tailed classification trained with all encountered data as the approximated upper bound.
Experimental settings. The experiments are conducted with NVIDIA GPU RTX8000 on CUDA 11.0. We also randomly split each dataset into multiple sessions. We run each algorithm ten times for each dataset and report their mean accuracy. We adopt ResNet18 [27] as the backbone network. For data augmentation, we use standard random crop and horizontal flips. In the base session training stage, we select top-$c\%$ weights at each layer and acquire the optimal soft-subnetworks with the best validation accuracy. In each incremental few-shot learning session, the total number of training epochs is $6$, and the learning rate is $0.02$. We train new class session samples using a few minor weights of the soft-subnetwork (Conv4x layer of ResNet18) obtained by the base session learning.
5 Results of Task-Incremental Learning
Table I: Performance comparisons of the proposed method and other state-of-the-art including baselines - PackNet [53] and SupSup [89] - on various benchmark datasets. We report the mean and standard deviation of the average accuracy (ACC) and average backward transfer (BWT) across $5$ independent runs with five seeds under the same experimental setup [17]. The best results are highlighted in bold. Also, $\dagger$    denotes results reported from [17].
Method CIFAR-100 Split CIFAR-100 Superclass TinyImageNet
ACC (%) BWT (%) ACC (%) BWT (%) ACC (%) BWT (%)
La-MaML [22] 71.37 ($\pm$ ${}^{\dagger}$5.39 ($\pm$ ${}^{\dagger}$ 54.44 ($\pm$ ${}^{\dagger}$6.65 ($\pm$ ${}^{\dagger}$ 66.90 ($\pm$ ${}^{\dagger}$9.13 ($\pm$ ${}^{\dagger}$
GPM [67] 73.18 ($\pm$ ${}^{\dagger}$1.17 ($\pm$ ${}^{\dagger}$ 57.33 ($\pm$ ${}^{\dagger}$0.37 ($\pm$ ${}^{\dagger}$ 67.39 ($\pm$ ${}^{\dagger}$  1.45 ($\pm$ ${}^{\dagger}$
FS-DGPM [17] 74.33 ($\pm$ ${}^{\dagger}$2.71 ($\pm$ ${}^{\dagger}$ 58.81 ($\pm$ ${}^{\dagger}$2.97 ($\pm$ ${}^{\dagger}$ 70.41 ($\pm$ ${}^{\dagger}$2.11 ($\pm$ ${}^{\dagger}$
PackNet [53] 72.39 ($\pm$ 0.4) 0.0 58.78 ($\pm$ 0.5) 0.0 55.46 ($\pm$ 1.2) 0.0
SupSup [89] 75.47 ($\pm$ 0.3) 0.0 61.70 ($\pm$ 0.3) 0.0 59.60 ($\pm$ 1.1) 0.0
${}^{\ast}$, $c=50\%$ 77.67 ($\pm$ 0.1) 0.0 61.58 ($\pm$ 0.0) 0.0 69.88 ($\pm$ 1.7) 0.0
WSN,   $c=50\%$ + FSO 79.00 ($\pm$ 0.3) 0.0 61.70 ($\pm$ 0.2) 0.0 72.04 ($\pm$ 0.7) 0.0
Multitask   79.75 ($\pm$ ${}^{\dagger}$   61.00 ($\pm$ ${}^{\dagger}$   77.10 ($\pm$ ${}^{\dagger}$
5.1 Comparisons with baselines in TIL
We use a multi-head setting to evaluate our WSN algorithm under the more challenging visual classification benchmarks. The WSN+FSO’s performances are compared with others in terms of two measurements on three major benchmark datasets as shown in Table I. Our WSN+FSO outperformed all state-of-the-art, achieving the best average accuracy of 79.00%, 61.70%, and 72.04%. WSN+FSO is also a forget-free model (BWT = ZERO), aligned with architecture-based models such as PackNet, SupSup, and WSN in these experiments. In addition, to show the effectiveness of the large single-scale task performance, we prepare WSN+FSO trained on the ImageNet-1K dataset, as shown in Table IX. WSN+FSO outperformed all baselines.
5.2 Statistics of WSN+FSO’s Representations in TIL
In the TinyImageNet task, the TinyImageNet takes 4 convolutional operators, followed by 3 fully connected layers. These operations represent more abstract features pooled by high-frequency components while losing low-frequency ones. To compensate for low-frequency components, we add an FSO to CNN architecture as shown in Figure 4. The lower layer’s output $\bm{x}^{l=2}$ is merged into the $l=3$th layer residual block through the FSO to acquire spatially ensembling features $\bm{x}^{l=3}$. The FSO also provides additional parameters to push the residual to zero. We will show the differences between ensembling features (WSN+FSO) and single features (WSN) represented by residual blocks as shown in Figure 5: WSN+FSO provides lower variances of feature maps and higher frequency components than WSN. The ensemble of representations led to better performances in TIL.
Table II: PSNR results with Fourier Subnueral Operator (FSO) layer ($f$-NeRVnormal-$\ast$) on UVG17 Video Sessions with average PSNR and Backward Transfer (BWT) of PSNR. Note that $\ast$ denotes our reproduced results.
Method Video Sessions Avg. MS-SSIM BWT
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17
STL, NeRV [9] 39.63 - 36.06 - 37.35 - 41.23 - 38.14 - 31.86 - 37.22 - 32.45 - - - / -
STL, ${}^{\ast}$ 39.66 44.89 36.28 41.13 38.14 31.53 42.03 34.74 36.58 36.85 29.22 31.81 37.27 34.18 31.45 38.41 43.86 36.94 / -
EWC [${}^{\ast}$ 11.15 9.21 12.71 11.40 15.58 9.25 7.06 12.96 6.34 10.31 9.55 13.39 5.76 8.67 10.93 10.92 28.29 11.38 / -16.13
iCaRL [${}^{\ast}$ 24.31 28.25 22.19 22.74 22.84 16.55 29.37 17.92 16.65 27.43 13.64 16.42 24.02 21.60 19.40 18.60 26.46 21.67 /   -6.23
ESMER [${}^{\ast}$ 30.77 26.33 22.79 21.35 23.76 13.64 28.25 15.22 16.71 23.78 13.35 15.23 18.21 19.22 24.59 20.61 22.42 20.95 / -15.23
${}^{\ast}$, c = 30.0 % 31.50 34.37 31.00 32.38 29.26 23.08 31.96 22.64 22.07 33.48 18.34 20.45 27.21 24.33 23.09 21.23 29.13 26.80 /  0.0
${}^{\ast}$, c = 50.0 % 34.02 34.93 31.04 31.74 28.95 23.07 31.26 22.32 21.93 33.35 18.22 20.34 26.88 24.22 22.72 21.30 28.86 26.77 /  0.0
WSN, c = 30.0 % + $f$-NeRV2 32.01 35.84 32.97 35.17 31.24 24.82 36.01 25.85 24.83 35.76 20.50 22.79 30.40 27.37 25.52 25.40 32.70 29.36 /  0.0
WSN, c = 30.0 % + $f$-NeRV3 33.64 39.24 34.21 37.79 34.05 27.17 38.17 29.79 26.56 36.18 22.97 24.36 32.50 30.22 27.62 29.15 35.68 31.72 /  0.0
MLT (upper-bound) 32.39 34.35 31.45 34.03 30.70 24.53 37.13 27.83 23.80 34.69 20.77 22.37 32.71 28.00 25.89 26.40 33.16 29.42 / -
t=0 t=1 t=2 t=3
Refer to caption Refer to caption Refer to caption Refer to caption
WSN (31.00, PSNR)
Refer to caption Refer to caption Refer to caption Refer to caption
WSN, $f$-NeRV3 (34.21, PSNR)
Refer to caption Refer to caption Refer to caption Refer to caption
WSN (29.26, PSNR)
Refer to caption Refer to caption Refer to caption Refer to caption
WSN, $f$-NeRV3 (34.05, PSNR)
Figure 6: Video Generation (from t=0 to t=3) with $c=30.0\%$ on the UVG17 dataset.
6 Results of Video Incremental Learning
6.1 Comparisons with Baselines
Peak signal-to-noise ratio (PSNR). To compare WSN+FSO with conventional representative continual learning methods such as EWC, iCaRL, ESMER, and WSN+FSO, we prepare the reproduced results, as shown in Table II. The architecture-based WSN outperformed the regularized method and replay method. The sparseness of WSN does not significantly affect sequential video representation results on two sequential benchmark datasets. Our WSN+FSO outperforms all conventional baselines including WSN and MLT (upper-bound of WSN) on the UVG17 benchmark datasets. Moreover, our performances of WSN with $f$-NeRV3 are better than those of $f$-NeRV2 since $f$-NeRV3 tends to represent local textures, stated in the following Section 6.2. Note that the number of parameters of MLT is precisely the same as those of WSN.
Compression. We follow NeRV’s video quantization and compression pipeline [10], except for the model pruning step, to evaluate performance drops and backward transfer in the video sequential learning, as shown in Figure 10. Once sequential training is done, our WSN+FSO doesn’t need any extra pruning and finetuning steps, unlike NeRV. This point is our key advantage of WSN+FSO over NeRV. Figure 10 (a) shows the results of various sparsity and bit-quantization on the UVG17 datasets: the 8bit WSN+FSO’s performances are comparable with 32bit ones without a significant video quality drop. From our observations, the 8-bit subnetwork seems to be enough for video implicit representation. Figure 10 (b) shows the rate-distortion curves. We compare WSN+FSO with WSN and NeRV (STL). For a fair comparison, we take steps of pruning, fine-tuning, quantizing, and encoding NeRV. Our WSN+FSO outperforms all baselines.
Performance and Capacity. Our WSN+FSO outperforms WSN and MTL, as stated in Figure 11 (a). This result might suggest that properly selected weights in Fourier space lead to generalization more than others in VIL. Moreover, to show the behavior of FSO, We prepare a progressive WSN+FSO’s capacity and investigate how FSO reuses weights over sequential video sessions, as shown in Figure 11 (b). WSN+FSO tends to progressively transfer weights used for a prior session to weights for new ones, but the proposition of reused weights gets smaller as video sessions increase.
6.2 WSN+FSO’s Video Representations
We prepare the results of video generation as shown in Figure 6. We demonstrate that a sparse solution (WSN with $c=30.0\%$, $f$-NeRV3) generates video representations sequentially without significant performance drops. Compared with WSN, WSN+FSO provides more precise representations. To find out the results, we inspect the layer-wise representations as shown in Figure 7, which offers essential observations that WSN+FSO tends to capture local textures broadly at the NeRV3 layer while WSN focuses on local objects. This WSN+FSO behavior could lead to more generalized performances. Moreover, we conduct an ablation study to inspect the sparsity-wise performances of $f$-NeRV3 while holding the remaining parameters’ sparsity (c=50.0 %), as shown in Figure 8. We could observe that as the sparsity of $f$-NeRV3 increases, its performances drop. This leads us to the importance of $f$-NeRV3’s representations.
NeRV3 NeRV4 NeRV5 NeRV6 Head
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
WSN, c=50.0% (28.95, PSNR)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
WSN, c=50.0%, $f$-NeRV2 (31.24, PSNR)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
WSN, c=50.0%, $f$-NeRV3 (34.05, PSNR)
Figure 7: WSN’s Representations of NeRV Blocks with $c=50.0\%$ on the UVG17 dataset.
NeRV3 NeRV4 NeRV5 NeRV6 Head
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PFNR, c=50.0%, sparsity of $f$-NeRV3=0.5(31.06, PSNR)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
WSM, c=50.0%, sparsity of $f$-NeRV3=5.0(32.38, PSNR)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
WSN, c=50.0%, sparsity of $f$-NeRV3=50.0(34.05, PSNR)
Figure 8: Various sparsity of $f$-NeRV3 ranging from 0.05 % (top row) to 50.0 % (bottom row) on the UVG17 dataset.
Refer to caption Refer to caption
(a) feature map variance (b) high-freq. of feature map
Figure 9: The comparisons of WSN of NeRV Blocks (blue area) with FSO (white area) in terms of Feature variances and high-frequency components: the (a) offers the variance of the feature map and the (b) provides $\Delta\log$ amplitudes at high-frequency ($1.0\pi$).
Refer to caption Refer to caption
(a) Quan. and Comp. (b) PSNR v.s. BPP
Figure 10: PSNR v.s. Bits-per-pixel (BPP) on the UVG17 datasets.
Refer to caption Refer to caption
(a) PSNR, $c=30.0\%$ (b) Capacity, $c=30.0\%$
Figure 11: WSN+FSO’s Comparison of PSNR with others and layer-wise accumulated capacities on the UVG17 dataset. Note that, in (b), green represents the percentage of reused subnetwork’s parameters of Stem, $f$-NeRV3, and NeRV5 at the current session (s) obtained at the past (s-1) video sessions
6.3 Statistics of WSN+FSO’s Representations in VIL
We provide the statistics of WSN+FSO’s video representations, as shown in Figure 9. In the video incremental task, the NeRV architecture takes stems of 2 fully connected layers, followed by 5 convolutional operators. These operations represent an image up-scaled by multiple scalars while losing high-frequency representations (see WSN’s representations of Figure 7). To compensate for high-frequency components, we add an FSO to NeRV blocks as shown in Figure 3. The Conv. layer’s output is merged into the outputs of FSO to acquire spatially ensembling features. We will show the differences between ensembling features (WSN+FSO) and single features (WSN), as shown in Figure 9: WSN+FSO provides a high variance of representations and higher frequency components than WSN at NeRV block 3. The ensemble of representations led to better performances in VIL, supporting the representations of Figure 7.
6.4 Ablation Studies of WSN+FSO
Variations of FSO. We prepared several ablation studies to prove the effectiveness of FSO. First, we show the performances of only real part (ignore an imaginary part) in f-NeRV2/3 as shown in Table III. The PSNR performances of only real part were lower than those of both real and imaginary parts in f-NeRV2/3. We infer that the imaginary part of the winning ticket improves the implicit neural representations. Second, we also investigate the effectiveness of only FSO without Conv. Layer in f-NeRV2/3, as shown in Table IV. The PSNR performances were lower than FSO with Conv block. Therefore, the ensemble of FSO and Conv improves the implicit representations. Lastly, we investigate the effectiveness of sparse FSO in STL, as shown in Table V. The sparse FSO boots the PSNR performances in STL. These ablation studies further strengthen the effectiveness of FSO for sequential neural implicit representations.
[…]
Method Video Sessions Avg. PSNR / BWT
1 2 3 4 5 6 7 8
$f$-NeRV2 34.46 33.91 32.17 36.43 25.26 20.74 30.18 25.45 29.82 / 0.0
$f$-NeRV2 w/o imag. 34.34 33.79 32.04 36.40 25.11 20.59 30.17 25.27 29.71 / 0.0
$f$-NeRV3 36.45 35.15 35.10 38.57 28.07 23.06 32.83 27.70 32.12 / 0.0
$f$-NeRV3 w/o imag. 35.66 34.65 34.09 37.95 25.80 21.94 32.17 26.91 31.15 / 0.0
[…]
Method Video Sessions Avg. PSNR / BWT
1 2 3 4 5 6 7 8
$f$-NeRV2 34.46 33.91 32.17 36.43 25.26 20.74 30.18 25.45 29.82 / 0.0
$f$-NeRV2 w/o conv. 30.05 32.10 30.12 31.82 24.00 19.60 28.21 24.47 27.54 / 0.0
$f$-NeRV3 36.45 35.15 35.10 38.57 28.07 23.06 32.83 27.70 32.12 / 0.0
$f$-NeRV3 w/o conv. 35.46 35.06 34.98 38.23 28.00 22.98 32.57 27.45 31.84 / 0.0
Table V: WSN+Fourier Subnueral Operator (FSO) layer ($f$-NeRVnormal-$\ast$) on UVG8 Video Sessions with average PSNR and Backward Transfer (BWT).
Method Video Sessions Avg. PSNR / BWT
1 2 3 4 5 6 7 8
STL, NeRV [${}^{\ast}$ 39.66 36.28 38.14 42.03 36.58 29.22 37.27 31.45 36.33 /   -
STL, NeRV , $f$-NeRV2 39.73 36.30 38.29 42.03 36.64 29.25 37.35 31.65 36.40 /   -
STL, NeRV , $f$-NeRV3 42.75 37.65 42.05 42.36 40.01 34.21 40.15 36.15 39.41 /   -
Forget-free Transfer Matrix. We prepare the transfer matrix to prove our WSN+FSO’s forget-freeness and to show video correlation among other videos, as shown in Figure 12 on the UVG17 dataset; lower triangular estimated by each session subnetwork denotes that our WSN+FSO is a forget-free method and upper triangular calculated by current session subnetwork denotes the video similarity between source and target. The WSN+FSO proves the effectiveness from the lower triangular of Figure 12 (a) and (b). Nothing special is observable from the upper triangular since they are not correlated, however, there might be some shared representations.
Refer to caption Refer to caption
(a) WSN, $c=30.0\%$ (b) WSN, $c=30.0\%$, $f$-NeRV3
Figure 12: Transfer Matrixes on the UVG17 dataset measured by PSNR of source and target videos.
7 Results On Few-shot CIL (FSCIL)
7.1 Results and Comparisons
Table VI: Classification accuracy of ResNet18 on CIFAR-100 for 5-way 5-shot incremental learning with the same class split as in TOPIC [14]. ${}^{\ast}$ denotes the results reported from [71]. ${}^{\dagger}$ represents our reproduced results.
Method sessions The gap with cRT
1 2 3 4 5 6 7 8 9
cRT [${}^{\ast}$ 72.28 69.58 65.16 61.41 58.83 55.87 53.28 51.38 49.51
TOPIC [14] 64.10 55.88 47.07 45.16 40.11 36.38 33.96 31.55 29.37 -20.14
CEC [97] 73.07 68.88 65.26 61.19 58.09 55.57 53.22 51.34 49.14 -0.37
F2M [71] 71.45 68.10 64.43 60.80 57.76 55.26 53.53 51.57 49.35 -0.16
LIMIT [99] 73.81 72.09 67.87 63.89 60.70 57.77 55.67 53.52 51.23 +1.72
MetaFSCIL [15] 74.50 70.10 66.84 62.77 59.48 56.52 54.36 52.56 49.97 +0.46
ALICE [62] 79.00 70.50 67.10 63.40 61.20 59.20 58.10 56.30 54.10 +4.59
Entropy-Reg [49] 74.40 70.20 66.54 62.51 59.71 56.58 54.52 52.39 50.14 +0.63
C-FSCIL [28] 77.50 72.45 67.94 63.80 60.24 57.34 54.61 52.41 50.23 +0.72
FSLL [54] 64.10 55.85 51.71 48.59 45.34 43.25 41.52 39.81 38.16 -11.35
FACT [98] 74.60 72.09 67.56 63.52 61.38 58.36 56.28 54.24 52.10 +2.59
WaRP  [36] 80.31 75.86 71.87 67.58 64.39 61.34 59.15 57.10 54.74 +5.23
SoftNet,   $c=90\%$ 79.97 75.75 71.76 67.36 64.09 60.91 59.07 56.94 54.76 +5.25
SoftNet,   $c=90\%$ + FSO 80.40 76.06 72.43 68.43 65.54 62.27 60.13 58.15 56.00 +6.49
Table VII: Classification accuracy of ResNet18 on miniImageNet for 5-way 5-shot incremental learning with the same class split as in TOPIC [14]. ${}^{\ast}$ denotes results reported from [71].
Method sessions The gap with cRT
1 2 3 4 5 6 7 8 9
cRT [${}^{\ast}$ 72.08 68.15 63.06 61.12 56.57 54.47 51.81 49.86 48.31 -
TOPIC [14] 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42 -23.89
IDLVQ-C [11] 64.77 59.87 55.93 52.62 49.88 47.55 44.83 43.14 41.84 -6.47
CEC [97] 72.00 66.83 62.97 59.43 56.70 53.73 51.19 49.24 47.63 -0.68
F2M [71] 72.05 67.47 63.16 59.70 56.71 53.77 51.11 49.21 47.84 -0.43
LIMIT [99] 73.81 72.09 67.87 63.89 60.70 57.77 55.67 53.52 51.23 +2.92
MetaFSCIL [15] 72.04 67.94 63.77 60.29 57.58 55.16 52.90 50.79 49.19 +0.88
ALICE [62] 80.60 70.60 67.40 64.50 62.50 60.00 57.80 56.80 55.70 +7.39
C-FSCIL [28] 76.40 71.14 66.46 63.29 60.42 57.46 54.78 53.11 51.41 +3.10
Entropy-Reg [49] 71.84 67.12 63.21 59.77 57.01 53.95 51.55 49.52 48.21 -0.10
Subspace Reg. [1] 80.37 71.69 66.94 62.53 58.90 55.00 51.94 49.76 46.79 -1.52
FSLL [54] 66.48 61.75 58.16 54.16 51.10 48.53 46.54 44.20 42.28 -6.03
FACT [98] 72.56 69.63 66.38 62.77 60.60 57.33 54.34 52.16 50.49 +2.18
WaRP [36] 72.99 68.10 64.31 61.30 58.64 56.08 53.40 51.72 50.65 +2.34
SoftNet,   $c=85\%$ 79.50 74.54 70.29 66.39 63.35 60.38 57.32 55.22 53.92 +5.61
SoftNet,   $c=85\%$ + FSO 79.72 74.72 70.73 66.88 64.05 61.82 58.03 56.01 54.80 +6.49
Table VIII: Classification accuracy of ResNet18 on CUB-200-2011 for 10-way 5-shot incremental learning (TOPIC class split [81]). ${}^{\ast}$ denotes results reported from [71]. ${}^{\dagger}$ represents our reproduced results.
Method sessions The gap with cRT
1 2 3 4 5 6 7 8 9 10 11
cRT [${}^{\ast}$ 77.16 74.41 71.31 68.08 65.57 63.08 62.44 61.29 60.12 59.85 59.30 -
TOPIC [14] 68.68 62.49 54.81 49.99 45.25 41.40 38.35 35.36 32.22 28.31 26.28 -34.80
SPPR [100] 68.68 61.85 57.43 52.68 50.19 46.88 44.65 43.07 40.17 39.63 37.33 -21.97
CEC [97] 75.85 71.94 68.50 63.50 62.43 58.27 57.73 55.81 54.83 53.52 52.28 -7.02
F2M [71] 77.13 73.92 70.27 66.37 64.34 61.69 60.52 59.38 57.15 56.94 55.89 -3.41
LIMIT [99] 75.89 73.55 71.99 68.14 67.42 63.61 62.40 61.35 59.91 58.66 57.41 -1.89
MetaFSCIL [15] 75.90 72.41 68.78 64.78 62.96 59.99 58.30 56.85 54.78 53.82 52.64 -6.66
ALICE [62] 77.40 72.70 70.60 67.20 65.90 63.40 62.90 61.90 60.50 60.60 60.10 -0.02
Entropy-Reg [49] 75.90 72.14 68.64 63.76 62.58 59.11 57.82 55.89 54.92 53.58 52.39 -6.91
FSLL [54] 72.77 69.33 65.51 62.66 61.10 58.65 57.78 57.26 55.59 55.39 54.21 -6.87
FACT [98] 75.90 73.23 70.84 66.13 65.56 62.15 61.74 59.83 58.41 57.89 56.94 -2.36
WaRP [36] 77.74 74.15 70.82 66.90 65.01 62.64 61.40 59.86 57.95 57.77 57.01 -2.29
SoftNet,  $c=90\%$ 78.07 74.58 71.37 67.54 65.37 62.60 61.07 59.37 57.53 57.21 56.75 -2.55
SoftNet,  $c=90\%$ + FSO 78.24 74.73 71.37 67.54 65.54 62.80 61.92 59.54 57.86 57.72 56.84 -2.46
Comparisons with SOTA. We compare SoftNet+FSO with the following state-of-art-methods on TOPIC class split [81] of three benchmark datasets - CIFAR100 (Table VI), miniImageNet (Table VII), and CUB-200-2011 (Table VIII). Leveraged by regularized backbone ResNet, SoftNet+FSO outperformed all existing architecture baselines FSLL [54], FACT [98], and WaRP [36] on CIFAR100, miniImageNet. On CUB-200-201, the performances of SoftNet+FSO are comparable with those of ALICE and LIMIT, considering that ALICE used class/data augmentations and LIMIT added an extra multi-head attention layer. Before finetuning SoftNet+FSO on CUB-200-2011, we pre-trained WSN+FSO on the ImageNet-1K dataset, as shown in Table IX. The WSN+FSO outperforms WSN and baselines.
Table IX: Image Classification Performances on ImageNet-1K.
Method Acc@1 Acc@5
ResNet18 [26] 69.75 89.07
WSN, c=99.0 % 69.46 89.05
WSN, c=99.0 % + FSO 70.63 89.84
8 Conclusion
Inspired by Regularized Lottery Ticket Hypothesis (RLTH), which states that competitive subnetworks exist within a dense network in continual learning tasks, we introduce an interpretable continual learning approach referred to as Winning Subnetworks, WSN, which leverages re-used weights within dense networks in Task Incremental Learning (TIL) scenarios. We have also introduced variants of WSN, such as Soft-subnetwork (SoftNet), to address the overfitting in few-shot class incremental learning (FSCIL). Additionally, To overcome the limitation of WSN (sparse re-used weights) in video incremental learning (VIL), we have a newly proposed module as another variant of WSN that aims to find an adaptive and compact sub-module referred to as Fourier Subneural Operator (FSO) in Fourier space to encode videos in each video training session. The FSO finds reusable winning subnetworks in Fourier space, providing various bandwidths. We extend Fourier representations to various continual learning scenarios such as VIL, TIL, and FSCIL. Extensive experiments demonstrate the effectiveness of FSO in three continual learning scenarios. Overall, FSO’s representation markedly enhanced task performance at different convolutional representational levels, the higher layers for TIL and FSCIL and the lower layers for VIL.
Acknowledgement.This work was supported by Institute for Information & communications Technology Promotion (IITP) grant funded by No. 2021-0-01381 and 2022-0-00184.